OpenStudy (anonymous):

how to write x-2y=12 in slope- intercept form

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

this pellet sucks it doesnt even help you.

jimthompson5910 (jim_thompson5910):

well I will help you, but you have to tell me how far you got

jimthompson5910 (jim_thompson5910):

ok I'll show you have to do this, but in the future please post your work so far

OpenStudy (anonymous):

Help meeeee ! i dont even know how to use this thing -_-

jimthompson5910 (jim_thompson5910):

\[\large x-2y=12 \] \[\large -2y=12-x \] \[\large -2y=-x+12 \] \[\large y=\frac{-x+12}{-2} \] \[\large y=\frac{-x}{-2}+\frac{12}{-2} \] \[\large y=\frac{1}{2}x-6 \]

jimthompson5910 (jim_thompson5910):

The idea is to isolate y and get it into y = mx+b form. The work above shows you how to do this.

OpenStudy (anonymous):

Thanks. How do you do x+2y-2=0

jimthompson5910 (jim_thompson5910):

x+2y-2=0 x+2y-2+2=0+2 x+2y = 2 2y = 2 - x I'll let you finish up

OpenStudy (anonymous):

do i just write it out like that ? I have to show work

jimthompson5910 (jim_thompson5910):

yes I'm showing the work (not the full work since you need to finish up)

OpenStudy (anonymous):

why cant you just finish it up ?

jimthompson5910 (jim_thompson5910):

There are only a few steps left and you should get used to doing your own work.

OpenStudy (anonymous):

I do but i just need to hurry up and do i need to turn it in.

jimthompson5910 (jim_thompson5910):

your next step is to divide both sides by 2 (since you have 2y and you want to isolate y) what do you get when you do that?

OpenStudy (anonymous):

y= 1 over 2 ?

jimthompson5910 (jim_thompson5910):

\[\large x+2y-2=0\] \[\large x+2y-2+2=0+2\] \[\large x+2y = 2\] \[\large 2y = 2 - x\] \[\large 2y = -x+2\] \[\large y = \frac{-x+2}{2}\] What's next?

OpenStudy (anonymous):

idk. But gtg thanks for your help